# Stochastic Gradient Langevin Unlearning

### v1

@@ Appendix C Existence of limiting distribution
We follow a similar analysis as in Lemma 3.5 but with two key differences: 1) our initial $W_{\infty}$ distance is not zero and 2) we are applying the same $c$-CNI. Let us slightly abuse the notation to denote the process $x_{t}^{j,\prime}\sim\nu_{t}^{j,\prime}$ be the process of learning on $\mathcal{D}$ as well but starting with $\nu_{0}^{0,\prime}=\nu_{\mathcal{D}|\mathcal{B}}$. As before, we first establish one iteration bound with epoch $t$, then iterate over $n/b$ iterations, and then iterate over $T$ to complete the proof. Consider the two CNI processes $x_{t}^{j},x_{t}^{j,\prime}$ with the same update $\psi_{j}(x)=x-\frac{\eta}{b}\sum_{i\in\mathcal{B}^{j}}\nabla f(x;\mathbf{d}_{i})$, where $x_{0}^{0}\sim\nu_{0}^{0}$ and $x_{0}^{0,\prime}\sim\nu_{\mathcal{D}|\mathcal{B}}$. Then using the same coupling construction as in the proof of Lemma 3.5 (i.e., using the same $W_{t}^{j}$), we have
@@
where (a) is due to the fact that projection $\Pi_{\mathcal{C}_{R}}$ is $1$-contractive, (b) is due to our coupling choice on the noise distribution, and (c) is due to the fact that $\psi_{j}$ is $(1-\eta m)$ contractive and the coupling choice on the mini-batch. By taking the essential supremum and infimum over all coupling over $\nu_{t|\mathcal{B}}^{j},\nu_{t|\mathcal{B}}^{j,\prime}$, we have
$W_{\infty}(\nu_{t|\mathcal{B}}^{j+1},\nu_{t}^{j+1|\mathcal{B},\prime})\leq(1-\eta m)W_{\infty}(\nu_{t|\mathcal{B}}^{j},\nu_{t|\mathcal{B}}^{j,\prime}).$ (63)
Iterating this bound gives a per epoch bound as follows.
$W_{\infty}(\nu_{t+1|\mathcal{B}}^{0},\nu_{t+1|\mathcal{B}}^{0,\prime})\leq(1-\eta m)^{n/b}W_{\infty}(\nu_{t|\mathcal{B}}^{0},\nu_{t|\mathcal{B}}^{0,\prime}).$ (64)
Note that by stationarity of $\nu_{\mathcal{D}|\mathcal{B}}$, we know that $\nu_{t|\mathcal{B}}^{0,\prime}=\nu_{\mathcal{D}|\mathcal{B}}$ for all $t$. Thus we have
$W_{\infty}(\nu_{t+1|\mathcal{B}}^{0},\nu_{\mathcal{D}|\mathcal{B}})\leq(1-\eta m)^{n/b}W_{\infty}(\nu_{t|\mathcal{B}}^{0},\nu_{\mathcal{D}|\mathcal{B}}).$ (65)
Finally, iterate this bound over $T$ epochs, and we complete the proof. ∎
Appendix K Experiment Details
K.1 $(\alpha,\varepsilon)$-RU to $(\epsilon,\delta)$-Unlearning Conversion
Let us first state the definition of $(\epsilon,\delta)$-unlearning from prior literature (Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021).
Definition K.1.
@@
Following the same proof of RDP-DP conversion (Proposition 3 in (Mironov, 2017)), we have the following $(\alpha,\varepsilon)$-RU to $(\epsilon,\delta)$-unlearning conversion as well.
Proposition K.2.
@@
K.2 Datasets
MNIST (Deng, 2012) contains the grey-scale image of number 00 to number $9$, each with $28\times 28$ pixels. We follow (Neel et al., 2021) to take the images with the label $3$ and $8$ as the two classes for logistic regression. The training data contains $11264$ instances in total and the testing data contains $1984$ samples. We spread the image into an $x\in\mathbb{R}^{d},d=724$ feature as the input of logistic regression.
CIFAR-10 (Krizhevsky et al., 2009) contains the RGB-scale image of ten classes for image classification, each with $32\times 32$ pixels. We also select class #3 (cat) and class #8 (ship) as the two classes for logistic regression. The training data contains $9728$ instances and the testing data contains $2000$ samples. We apply data pre-processing on CIFAR-10 by extracting the compact feature encoding from the last layer before pooling of an off-the-shelf pre-trained ResNet18 model (He et al., 2016) from Torch-vision library (maintainers & contributors, 2016; Paszke et al., 2019) as the input of our logistic regression. The compact feature encoding is $x\in\mathbb{R}^{d},d=512$.
All the inputs from the datasets are normalized with the $\ell_{2}$ norm of $1$. Note that We drop some date points compared with  (Chien et al., 2024) to make the number of training data an integer multiple of the maximum batch size in our experiment, which is 512.
K.3 Experiment Settings
@@
where the objective consists of a standard logistic regression loss $l(\textbf{w}^{\top}x_{i},y_{i})=-\log\sigma(y_{i}\textbf{w}^{\top}\textbf{x}_{i})$, where $\sigma(t)=\frac{1}{1+\exp(-t)}$ is the sigmoid function; and a $\ell_{2}$ regularization term where $\lambda$ is a hyperparameter to control the regularization, and we set $\lambda$ as $10^{-6}\times n$ across all the experiments. By simple algebra one can show that (Guo et al., 2020)
@@
where $\Pi_{\mathcal{C}_{M}}$ denotes the gradient clipping projection into the Euclidean ball with the radius of $M$, to satisfy the Lipschitz constant bound. According to Proposition 5.2 of (Ye & Shokri, 2022), the per-sampling clipping operation still results in a $L$-smooth, $m$-strongly convex objective. The resulting stochastic gradient update on the full dataset is as follows:
@@
Learning from scratch set-up For the baselines and our stochastic gradient Langevin unlearning framework, we all sample the initial weight w randomly sampled from i.i.d Gaussian distribution $\mathcal{N}(\mu_{0},\frac{2\sigma^{2}}{m})$, where $\mu_{0}$ is a hyper-parameter denoting the initialization mean and we set as 00. For the stochastic gradient langevin unlearning method, the burn-in steps $T$ w.r.t. different batch sizes are listed in Table. 2. we follow  (Chien et al., 2024) and set $T=10,000$ for the baselines (D2D and Langevin unlearning) to converge.
Table 2: The Burn-in step $T$ set for different batch sizes for the stochastic gradient Langevin unlearning framework
@@
Unlearning request implementation. In our experiment, for an unlearning request of removing data point $i$, we replace its feature with random features drawn from $\mathcal{N}(0,I_{d})$ and its label with a random label drawn uniformly at random drawn from all possible classes. This is similar to the DP replacement definition defined in (Kairouz et al., 2021), where they replace a point with a special null point $\perp$.
@@
$\bullet$ For the experiments in Fig. 3a, we calculate the noise to add after gradient descent with the non-private bound as illustrated in Theorem. L.1 (Theorem 9 in (Neel et al., 2021)); For experiments with sequential unlearning requests in Fig. 3b, we calculate the least step number and corresponding noise with the bound in Theorem. L.2(Theorem 28 in (Neel et al., 2021)).
$\bullet$ The implementation of D2D follows the pseudo code shown in Algorithm 1,2 in  (Neel et al., 2021) as follows:
@@
The settings and the calculation of $I,\sigma$ in Algorithm. 3 are discussed in the later part of this section and could be found in Section. L.
Langevin unlearning  (Chien et al., 2024)
We follow exactly the experiment details described in  (Chien et al., 2024).
General Implementation of Stochastic Gradient Langevin Unlearning (ours)
$\bullet$ We set the step size $\eta$ for the stochastic gradient Langevin unlearning framework across all the experiments as $1/L$.
$\bullet$ The pseudo-code for stochastic gradient Langevin unlearning framework is shown in Algorithm. 1.
K.4 Implementation Details for Fig. 3a
In this experiment, we first train the methods on the original dataset $\mathcal{D}$ from scratch to obtain the initial weights $\textbf{w}_{0}$. Then we randomly remove a single data point ($S=1$) from the dataset to get the new dataset $\mathcal{D}^{\prime}$, and unlearn the methods from the initial weights $\hat{\textbf{w}}$ and test the accuracy on the testing set. We follow  (Chien et al., 2024) and set the target $\hat{\epsilon}$ with $6$ different values as $[0.05,0.1,0.5,1,2,5]$. For each target $\hat{\epsilon}$:
$\bullet$ For D2D, we set two different unlearning gradient descent step budgets as $I=1,5$, and calculate the corresponding noise to be added to the weight after gradient descent on $\mathcal{D}$ according to Theorem. L.1.
$\bullet$ For the Langevin unlearning framework (Chien et al., 2024), we set the unlearning fine-tune step budget as $\hat{K}=1$ only, and calculate the smallest $\sigma$ that could satisfy the fine-tune step budget and target $\hat{\epsilon}$ at the same time. The calculation follows the binary search algorithm described in the original paper.
@@
Algorithm 4 Stochastic Gradient Langevin Unlearning: binary search $\sigma$ that satisfy $\hat{K}$ and target $\hat{\epsilon}$ budget
@@
Algorithm 5 Stochastic Gradient Langevin Unlearning [non-convergent]: find the least $K$ satisfying $\hat{\epsilon}$
@@
Table 3: $\sigma$ of stochastic gradient langevin unlearning.
@@
$\bullet$ For D2D, According to the original paper, only one data point could be removed at a time. We calculate the least required steps and the noise to be added according to Theorem. L.2.
$\bullet$ For Langevin unlearning, we fix the $\sigma=0.03$, and we let the model unlearn $[5,10,20]$ per time. The least required unlearning steps are obtained following Chien et al. (2024).
@@
Algorithm 6 Stochastic Gradient Langevin Unlearning: find the least unlearn step $K$ in sequential settings
@@
K.6 Implementation Details for Fig. 3c and 3d
In this study, we fix $S=100$ to remove, set different $\sigma=[0.05,0.1,0.2,0.5,1]$ and set batch size as $b=[32,128,512,\text{full batch}]$. We train the stochastic gradient langevin unlearning framework from scratch to get the initial weight, then remove some data, unlearn the model, and report the accuracy. We calculate the least required unlearning steps $K$ by calling Algorithm. 6.
Appendix L Unlearning Guarantee of Delete-to-Descent (Neel et al., 2021)
The discussion is similar to those in (Chien et al., 2024), we include them for completeness.
Theorem L.1 (Theorem 9 in (Neel et al., 2021), with internal non-private state).
Assume for all $\mathbf{d}\in\mathcal{X}$, $f(x;\mathbf{d})$ is $m$-strongly convex, $M$-Lipschitz and $L$-smooth in $x$. Define $\gamma=\frac{L-m}{L+m}$ and $\eta=\frac{2}{L+m}$. Let the learning iteration $T\geq I+\log(\frac{2Rmn}{2M})/\log(1/\gamma)$ for PGD (Algorithm 1 in (Neel et al., 2021)) and the unlearning algorithm (Algorithm 2 in (Neel et al., 2021), PGD fine-tuning on learned parameters before adding Gaussian noise) run with $I$ iterations. Assume $\epsilon=O(\log(1/\delta))$, let the standard deviation of the output perturbation gaussian noise $\sigma$ to be
@@
Theorem L.2 (Theorem 28 in (Neel et al., 2021), without internal non-private state).
Assume for all $\mathbf{d}\in\mathcal{X}$, $f(x;\mathbf{d})$ is $m$-strongly convex, $M$-Lipschitz and $L$-smooth in $x$. Define $\gamma=\frac{L-m}{L+m}$ and $\eta=\frac{2}{L+m}$. Let the learning iteration $T\geq I+\log(\frac{2Rmn}{2M})/\log(1/\gamma)$ for PGD (Algorithm 1 in (Neel et al., 2021)) and the unlearning algorithm (Algorithm 2 in (Neel et al., 2021), PGD fine-tuning on learned parameters after adding Gaussian noise) run with $I+\log(\log(4di/\delta))/\log(1/\gamma)$ iterations for the $i^{th}$ sequential unlearning request, where $I$ satisfies
@@
Note that the privacy guarantee of D2D (Neel et al., 2021) is with respect to add/remove dataset adjacency and ours is the replacement dataset adjacency. However, by a slight modification of the proof of Theorem L.1 and L.2, one can show that a similar (but slightly worse) bound of the theorem above also holds for D2D (Neel et al., 2021). For simplicity and fair comparison, we directly use the bound in Theorem L.1 and L.2 in our experiment. Note that (Kairouz et al., 2021) also compares a special replacement DP with standard add/remove DP, where a data point can only be replaced with a null element in their definition. In contrast, our replacement data adjacency allows arbitrary replacement which intuitively provides a stronger privacy notion.
The non-private internal state of D2D. There are two different versions of the D2D algorithm depending on whether one allows the server (model holder) to save and leverage the model parameter before adding Gaussian noise. The main difference between Theorem L.1 and L.2 is whether their unlearning process starts with the “clean” model parameter (Theorem L.1) or the noisy model parameter (Theorem L.2). Clearly, allowing the server to keep and leverage the non-private internal state provides a weaker notion of privacy (Neel et al., 2021). In contrast, our stochastic gradient Langevin unlearning approach by default only keeps the noisy parameter so that we do not save any non-private internal state. As a result, one should compare the stochastic gradient Langevin unlearning to D2D with Theorem L.2 for a fair comparison.
Appendix M Unlearning Guarantees of Langevin Unlearning (Chien et al., 2024)
In this section, we restate the main results of Langevin unlearning (Chien et al., 2024) and provide a detailed comparison and discussion for the case of strongly convex objective functions.
Theorem M.1 (RU guarantee of PNGD unlearning, strong convexity).
@@
Theorem M.2 (RDP guarantee of PNGD learning, strong convexity).
@@
On the other hand, combining our Theorem 3.2,  3.5 and using the worst case bound on $Z_{\mathcal{B}}$ along with some simplification, we have
@@
Now we compare the bound (79) obtained by Langevin dynamic in (Chien et al., 2024) and our bound (81) based on PABI analysis. First notice that the “initial distance” in (81) has an additional factor $1/(2m\eta)$. When we choose the largest possible step size $\eta=\frac{1}{L}$, then this factor is $\frac{L}{2m}$ which is most likely greater than $1$ in a real-world scenario unless the objective function is very close to a quadratic function. As a result, when $K$ is sufficiently small (i.e., $K=1$), possible that the bound (79) results in a better privacy guarantee when unlearning one data point. This is observed in our experiment, Figure 3a. Nevertheless, note that the decaying rate of (81) is independent of the Rényi divergence order $\alpha$, which actually leads to a much better rate in practice. More specifically, for the case $n=10^{4}$ (which is roughly our experiment setting). To achieve $(1,1/n)$-unlearning guarantee the corresponding $\alpha$ is roughly at scale $\alpha\approx 10$, since $\log(n)/(\alpha-1)$ needs to be less than $1$ in the $(\alpha,\varepsilon)$-RU to $(\epsilon,\delta)$-unlearning conversion (Proposition K.2). As a result, the decaying rate of (81) obtained by PABI analysis is superior by a factor of $2\alpha$, which implies a roughly $20$x saving in complexity for $K$ large enough.

### v2

@@ Appendix I Improved bound with randomized $\mathcal{B}$
We follow a similar analysis as in Lemma 3.3 but with two key differences: 1) our initial $W_{\infty}$ distance is not zero and 2) we are applying the same $c$-CNI. Let us slightly abuse the notation to denote the process $x_{t}^{j,\prime}\sim\nu_{t}^{j,\prime}$ be the process of learning on $\mathcal{D}$ as well but starting with $\nu_{0}^{0,\prime}=\nu_{\mathcal{D}|\mathcal{B}}$. As before, we first establish one iteration bound with epoch $t$, then iterate over $n/b$ iterations, and then iterate over $T$ to complete the proof. Consider the two CNI processes $x_{t}^{j},x_{t}^{j,\prime}$ with the same update $\psi_{j}(x)=x-\frac{\eta}{b}\sum_{i\in\mathcal{B}^{j}}\nabla f(x;\mathbf{d}_{i})$, where $x_{0}^{0}\sim\nu_{0}^{0}$ and $x_{0}^{0,\prime}\sim\nu_{\mathcal{D}|\mathcal{B}}$. We will use the similar coupling construction as in the proof of Lemma 3.3 (i.e., $W_{t}^{j}=W_{t}^{j,\prime}$). However, note that we will not restrict the coupling between the two initial distributions $\nu_{0}^{0}=\nu_{0|\mathcal{B}}$ and $\nu_{0}^{0,\prime}=\nu_{\mathcal{D}|\mathcal{B}}$. Let us denote the coupling for $W$ as $\gamma_{W}$ and the coupling for the initial distributions $\nu_{0}^{0}=\nu_{0|\mathcal{B}}$ and $\nu_{0}^{0,\prime}=\nu_{\mathcal{D}|\mathcal{B}}$ as $\gamma_{I}$. We denote the overall coupling as $\gamma=(\gamma_{I},\gamma_{W})$. For our specific choice of coupling $\gamma_{W}$, we have
@@
where (a) is due to the fact that projection $\Pi_{\mathcal{C}_{R}}$ is $1$-contractive, (b) is due to our coupling choice on the noise distribution, and (c) is due to the fact that $\psi_{j}$ is $(1-\eta m)$ contractive and the coupling choice on the mini-batch.
Now we iterate the bound above within the epoch and then over $T$ epoch similar as before, which leads to the per epoch bound as follows
$\text{esssup}_{\gamma}\|x_{T}^{0}-x_{T}^{0,\prime}\|\leq\text{esssup}_{\gamma}(1-\eta m)^{Tn/b}\|x_{0}^{0}-x_{0}^{0,\prime}\|=\text{esssup}_{\gamma_{I}}(1-\eta m)^{Tn/b}\|x_{0}^{0}-x_{0}^{0,\prime}\|.$ (82)
Note that this holds for all coupling $\gamma_{I}$ for initial distributions. Let us now choose a specific (tie break arbitrary) $\gamma_{I}$ so that the $W_{\infty}(\nu_{0|\mathcal{B}},\nu_{\mathcal{D}|\mathcal{B}})=\text{esssup}_{\gamma_{I}}\|x_{0}^{0}-x_{0}^{0,\prime}\|$ is attained. For this specific coupling $\gamma_{I}$, we have
$\text{esssup}_{(\gamma_{I},\gamma_{W})}\|x_{T}^{0}-x_{T}^{0,\prime}\|\leq(1-\eta m)^{Tn/b}W_{\infty}(\nu_{0|\mathcal{B}},\nu_{\mathcal{D}|\mathcal{B}}).$ (83)
Finally, by the definition of infimum and observe that $x_{T}^{0,\prime}\sim\nu_{\mathcal{D}|\mathcal{B}}$ due to the stationary property of $\nu_{\mathcal{D}|\mathcal{B}}$ (which is independent of our choice of coupling), we have
$W_{\infty}(\nu_{T|\mathcal{B}}^{0},\nu_{\mathcal{D}|\mathcal{B}})\leq\text{esssup}_{(\gamma_{I},\gamma_{W})}\|x_{T}^{0}-x_{T}^{0,\prime}\|\leq(1-\eta m)^{Tn/b}W_{\infty}(\nu_{0|\mathcal{B}},\nu_{\mathcal{D}|\mathcal{B}}).$ (84)
Hence we complete the proof. ∎
Appendix N Experiment Details
N.1 $(\alpha,\varepsilon)$-RU to $(\epsilon,\delta)$-Unlearning Conversion
Let us first state the definition of $(\epsilon,\delta)$-unlearning from prior literature [7, 8, 9].
Definition N.1.
@@
Following the same proof of RDP-DP conversion (Proposition 3 in [19]), we have the following $(\alpha,\varepsilon)$-RU to $(\epsilon,\delta)$-unlearning conversion as well.
Proposition N.2.
@@
N.2 Datasets
MNIST [25] contains the grey-scale image of number 00 to number $9$, each with $28\times 28$ pixels. We follow [9] to take the images with the label $3$ and $8$ as the two classes for logistic regression. The training data contains $11264$ instances in total and the testing data contains $1984$ samples. We spread the image into an $x\in\mathbb{R}^{d},d=724$ feature as the input of logistic regression.
CIFAR-10 [26] contains the RGB-scale image of ten classes for image classification, each with $32\times 32$ pixels. We also select class #3 (cat) and class #8 (ship) as the two classes for logistic regression. The training data contains $9728$ instances and the testing data contains $2000$ samples. We apply data pre-processing on CIFAR-10 by extracting the compact feature encoding from the last layer before pooling of an off-the-shelf pre-trained ResNet18 model [27] from Torch-vision library [34, 35] as the input of our logistic regression. The compact feature encoding is $x\in\mathbb{R}^{d},d=512$.
All the inputs from the datasets are normalized with the $\ell_{2}$ norm of $1$. Note that We drop some date points compared with  [11] to make the number of training data an integer multiple of the maximum batch size in our experiment, which is 512.
N.3 Experiment Settings
Hardware and Frameworks All the experiments run with PyTorch=2.1.2 [36] and numpy=1.24.3 [37]. The codes run on a server with a single NVIDIA RTX 6000 GPU with AMD EPYC 7763 64-Core Processor.
@@
where the objective consists of a standard logistic regression loss $l(\textbf{w}^{\top}x_{i},y_{i})=-\log\sigma(y_{i}\textbf{w}^{\top}\textbf{x}_{i})$, where $\sigma(t)=\frac{1}{1+\exp(-t)}$ is the sigmoid function; and a $\ell_{2}$ regularization term where $\lambda$ is a hyperparameter to control the regularization, and we set $\lambda$ as $10^{-6}\times n$ across all the experiments. By simple algebra one can show that [7]
@@
where $\Pi_{\mathcal{C}_{M}}$ denotes the gradient clipping projection into the Euclidean ball with the radius of $M$, to satisfy the Lipschitz constant bound. According to Proposition 5.2 of [22], the per-sampling clipping operation still results in a $L$-smooth, $m$-strongly convex objective. The resulting stochastic gradient update on the full dataset is as follows:
@@
Learning from scratch set-up For the baselines and our PNSGD unlearning framework, we all sample the initial weight w randomly sampled from i.i.d Gaussian distribution $\mathcal{N}(\mu_{0},\frac{2\sigma^{2}}{m})$, where $\mu_{0}$ is a hyper-parameter denoting the initialization mean and we set as 00. For the PNSGD unlearning method, the burn-in steps $T$ w.r.t. different batch sizes are listed in Table. 2. we follow  [11] and set $T=10,000$ for the baselines (D2D and Langevin unlearning) to converge.
Table 2: The Burn-in step $T$ set for different batch sizes for the PNSGD unlearning framework
@@
Unlearning request implementation. In our experiment, for an unlearning request of removing data point $i$, we replace its feature with random features drawn from $\mathcal{N}(0,I_{d})$ and its label with a random label drawn uniformly at random drawn from all possible classes. This is similar to the DP replacement definition defined in [38], where they replace a point with a special null point $\perp$.
@@
$\bullet$ For the experiments in Fig. 3a, we calculate the noise to add after gradient descent with the non-private bound as illustrated in Theorem. O.1 (Theorem 9 in [9]); For experiments with sequential unlearning requests in Fig. 3b, we calculate the least step number and corresponding noise with the bound in Theorem. O.2(Theorem 28 in [9]).
$\bullet$ The implementation of D2D follows the pseudo code shown in Algorithm 1,2 in  [9] as follows:
@@
The settings and the calculation of $I,\sigma$ in Algorithm. 3 are discussed in the later part of this section and could be found in Section. O.
Langevin unlearning  [11]
We follow exactly the experiment details described in  [11].
General Implementation of PNSGD Unlearning (ours)
$\bullet$ We set the step size $\eta$ for the PNSGD unlearning framework across all the experiments as $1/L$.
$\bullet$ The pseudo-code for PNSGD unlearning framework is shown in Algorithm. 1.
N.4 Implementation Details for Fig. 3a
In this experiment, we first train the methods on the original dataset $\mathcal{D}$ from scratch to obtain the initial weights $\textbf{w}_{0}$. Then we randomly remove a single data point ($S=1$) from the dataset to get the new dataset $\mathcal{D}^{\prime}$, and unlearn the methods from the initial weights $\hat{\textbf{w}}$ and test the accuracy on the testing set. We follow  [11] and set the target $\hat{\epsilon}$ with $6$ different values as $[0.05,0.1,0.5,1,2,5]$. For each target $\hat{\epsilon}$:
$\bullet$ For D2D, we set two different unlearning gradient descent step budgets as $I=1,5$, and calculate the corresponding noise to be added to the weight after gradient descent on $\mathcal{D}$ according to Theorem. O.1.
$\bullet$ For the Langevin unlearning framework [11], we set the unlearning fine-tune step budget as $\hat{K}=1$ only, and calculate the smallest $\sigma$ that could satisfy the fine-tune step budget and target $\hat{\epsilon}$ at the same time. The calculation follows the binary search algorithm described in the original paper.
@@
Algorithm 4 PNSGD Unlearning: binary search $\sigma$ that satisfy $\hat{K}$ and target $\hat{\epsilon}$ budget
@@
Algorithm 5 PNSGD Unlearning [non-convergent]: find the least $K$ satisfying $\hat{\epsilon}$
@@
Table 3: $\sigma$ of PNSGD unlearning.
@@
$\bullet$ For D2D, According to the original paper, only one data point could be removed at a time. We calculate the least required steps and the noise to be added according to Theorem. O.2.
$\bullet$ For Langevin unlearning, we fix the $\sigma=0.03$, and we let the model unlearn $[5,10,20]$ per time. The least required unlearning steps are obtained following [11].
@@
Algorithm 6 PNSGD Unlearning: find the least unlearn step $K$ in sequential settings
@@
N.6 Implementation Details for Fig. 3c and 3d
In this study, we fix $S=100$ to remove, set different $\sigma=[0.05,0.1,0.2,0.5,1]$ and set batch size as $b=[32,128,512,\text{full batch}]$. We train the PNSGD unlearning framework from scratch to get the initial weight, then remove some data, unlearn the model, and report the accuracy. We calculate the least required unlearning steps $K$ by calling Algorithm. 6.
Appendix O Unlearning Guarantee of Delete-to-Descent [9]
The discussion is similar to those in [11], we include them for completeness.
Theorem O.1 (Theorem 9 in [9], with internal non-private state).
Assume for all $\mathbf{d}\in\mathcal{X}$, $f(x;\mathbf{d})$ is $m$-strongly convex, $M$-Lipschitz and $L$-smooth in $x$. Define $\gamma=\frac{L-m}{L+m}$ and $\eta=\frac{2}{L+m}$. Let the learning iteration $T\geq I+\log(\frac{2Rmn}{2M})/\log(1/\gamma)$ for PGD (Algorithm 1 in [9]) and the unlearning algorithm (Algorithm 2 in [9], PGD fine-tuning on learned parameters before adding Gaussian noise) run with $I$ iterations. Assume $\epsilon=O(\log(1/\delta))$, let the standard deviation of the output perturbation gaussian noise $\sigma$ to be
@@
Theorem O.2 (Theorem 28 in [9], without internal non-private state).
Assume for all $\mathbf{d}\in\mathcal{X}$, $f(x;\mathbf{d})$ is $m$-strongly convex, $M$-Lipschitz and $L$-smooth in $x$. Define $\gamma=\frac{L-m}{L+m}$ and $\eta=\frac{2}{L+m}$. Let the learning iteration $T\geq I+\log(\frac{2Rmn}{2M})/\log(1/\gamma)$ for PGD (Algorithm 1 in [9]) and the unlearning algorithm (Algorithm 2 in [9], PGD fine-tuning on learned parameters after adding Gaussian noise) run with $I+\log(\log(4di/\delta))/\log(1/\gamma)$ iterations for the $i^{th}$ sequential unlearning request, where $I$ satisfies
@@
Note that the privacy guarantee of D2D [9] is with respect to add/remove dataset adjacency and ours is the replacement dataset adjacency. However, by a slight modification of the proof of Theorem O.1 and O.2, one can show that a similar (but slightly worse) bound of the theorem above also holds for D2D [9]. For simplicity and fair comparison, we directly use the bound in Theorem O.1 and O.2 in our experiment. Note that [38] also compares a special replacement DP with standard add/remove DP, where a data point can only be replaced with a null element in their definition. In contrast, our replacement data adjacency allows arbitrary replacement which intuitively provides a stronger privacy notion.
The non-private internal state of D2D. There are two different versions of the D2D algorithm depending on whether one allows the server (model holder) to save and leverage the model parameter before adding Gaussian noise. The main difference between Theorem O.1 and O.2 is whether their unlearning process starts with the “clean” model parameter (Theorem O.1) or the noisy model parameter (Theorem O.2). Clearly, allowing the server to keep and leverage the non-private internal state provides a weaker notion of privacy [9]. In contrast, our PNSGD unlearning approach by default only keeps the noisy parameter so that we do not save any non-private internal state. As a result, one should compare the PNSGD unlearning to D2D with Theorem O.2 for a fair comparison.
Appendix P Unlearning Guarantees of Langevin Unlearning [11]
In this section, we restate the main results of Langevin unlearning [11] and provide a detailed comparison and discussion for the case of strongly convex objective functions.
Theorem P.1 (RU guarantee of PNGD unlearning, strong convexity).
@@
Theorem P.2 (RDP guarantee of PNGD learning, strong convexity).
@@
On the other hand, combining our Theorem 3.2,  3.3 and using the worst case bound on $Z_{\mathcal{B}}$ along with some simplification, we have
@@
Now we compare the bound (98) obtained by Langevin dynamic in [11] and our bound (100) based on PABI analysis. First notice that the “initial distance” in (100) has an additional factor $1/(2m\eta)$. When we choose the largest possible step size $\eta=\frac{1}{L}$, then this factor is $\frac{L}{2m}$ which is most likely greater than $1$ in a real-world scenario unless the objective function is very close to a quadratic function. As a result, when $K$ is sufficiently small (i.e., $K=1$), possible that the bound (98) results in a better privacy guarantee when unlearning one data point. This is observed in our experiment, Figure 3a. Nevertheless, note that the decaying rate of (100) is independent of the Rényi divergence order $\alpha$, which actually leads to a much better rate in practice. More specifically, for the case $n=10^{4}$ (which is roughly our experiment setting). To achieve $(1,1/n)$-unlearning guarantee the corresponding $\alpha$ is roughly at scale $\alpha\approx 10$, since $\log(n)/(\alpha-1)$ needs to be less than $1$ in the $(\alpha,\varepsilon)$-RU to $(\epsilon,\delta)$-unlearning conversion (Proposition N.2). As a result, the decaying rate of (100) obtained by PABI analysis is superior by a factor of $2\alpha$, which implies a roughly $20$x saving in complexity for $K$ large enough.